OpenStudy (anonymous):

find the area of the region bounded by the line y=x+1 and curve y=x^2-1

OpenStudy (saifoo.khan):

Find the intersection points first.

OpenStudy (anonymous):

how do i do that? by setting the eqn to 0?

OpenStudy (saifoo.khan):

Solving them simultaneously.

OpenStudy (anonymous):

i got x=-1,1

OpenStudy (saifoo.khan):

x^2 -1 = x+1 x^2 -x -2 = 0

OpenStudy (saifoo.khan):

x = -1, x = 2

OpenStudy (anonymous):

thank you!

OpenStudy (saifoo.khan):

Now integrate with using limits.

OpenStudy (anonymous):

integrate what?

OpenStudy (saifoo.khan):

Here.